OpenStudy (anonymous):

An equation is shown below: (4x - 1) + (2x + 2) = 1 Which statement shows a correct next step in solving the equation? (1 point) Question 18 options: 1) The equation can become 6x + 1 = 1 by combining like terms. 2) The equation can become 3x + 4x = 1 by solving the terms in the brackets. 3) The equation can become 3x + 4x = 1 by applying the associative property of multiplication. 4) The equation can become 6x + 1 = 1 by applying the distributive property.

OpenStudy (anonymous):

(4x - 1) + (2x + 2) = 1 4x - 1 + 2x + 2 = 1 add

OpenStudy (anonymous):

6x +1 = 1

OpenStudy (anonymous):

The equation can become 6x + 1 = 1 by applying the distributive property.

OpenStudy (anonymous):

or The equation can become 6x + 1 = 1 by combining like terms.

OpenStudy (anonymous):

1 or 4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

what you think

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

1

OpenStudy (anonymous):

no 4

OpenStudy (anonymous):

1